OpenStudy (keke):

simplify the square root of 1,500

OpenStudy (anonymous):

38.73

OpenStudy (anonymous):

it is \[10\sqrt{15}\]

OpenStudy (angela210793):

20sqr15

OpenStudy (anonymous):

angela ????????????

OpenStudy (anonymous):

see 1500 = 15 x 10 x10 so u can take out a 10 n 15 is left

OpenStudy (anonymous):

5^3*2^3

OpenStudy (angela210793):

1500=25*60=25*4*15 oops....

OpenStudy (angela210793):

sorry....I've been studying math for 8 hrs with only 20 mins break :(:(:( I'm loosing my mind :(:(:(

OpenStudy (anonymous):

go take a walk or go to sleep ;)

OpenStudy (angela210793):

Can't..... have to study 1/2 yet :(:(:(

OpenStudy (anonymous):

then go offline and concentrate on your own studies....